In 12years Juliet will be 3 times as old as she was 8 years ago.How old is she now.

Answer by fractalier(6550)

Let J be her age NOW.
In 12 years she will be J + 12.
8 years ago she was J - 8.
Then
J + 12 = 3(J - 8)
J + 12 = 3J - 24
36 = 2J and
J = 18